\begin{thebibliography}{10}
  \bibitem{A1} Akhiezer, N. I.
\newblock Elements of the theory of elliptic functions.
\newblock Translated from the second Russian edition by H. H. McFaden. Translations of Mathematical Monographs, 79. American Mathematical Society, Providence, RI, 1990. viii+237 pp.

\bibitem{bargmann61}
V.~Bargmann.
\newblock On a {H}ilbert space of analytic functions and an associated integral
  transform.
\newblock {\em Comm. Pure Appl. Math.}, 14:187--214, 1961.

\bibitem{bargmann71}
V.~Bargmann, P.~Butera, L.~Girardello, and J.~R. Klauder.
\newblock On the completeness of coherent states.
\newblock {\em Rep. Math. Phys.}, 2:221--228, 1971.


% \bibitem{beur89}
% A.~Beurling.
% \newblock {\em The collected works of {A}rne {B}eurling. {V}ol. 2}.
% \newblock Contemporary Mathematicians. Birkh\"auser Boston Inc., Boston, MA,
%   1989.
% \newblock Harmonic analysis, Edited by L. Carleson, P. Malliavin, J. Neuberger
%   and J. Wermer.

\bibitem{FS17}
M.~Faulhuber and S.~Steinerberger.
\newblock Optimal {G}abor frame bounds for separable lattices and estimates for
  {J}acobi theta functions.
\newblock {\em J. Math. Anal. Appl.}, 445(1):407--422, 2017.

\bibitem{fg89jfa}
H.~G. Feichtinger and K.~Gr{\"o}chenig.
\newblock Banach spaces related to integrable group representations and their
  atomic decompositions. {I}.
\newblock {\em J. Functional Anal.}, 86(2):307--340, 1989.

\bibitem{feichtinger-kozek98}
H.~G. Feichtinger and W.~Kozek.
\newblock Quantization of {T}{F} lattice-invariant operators on elementary
  {L}{C}{A} groups.
\newblock In {\em Gabor analysis and algorithms}, pages 233--266. Birkh\"auser
  Boston, Boston, MA, 1998.

\bibitem{FZ98}
H.~G. Feichtinger and G.~Zimmermann.
\newblock A {B}anach space of test functions for {G}abor analysis.
\newblock In {\em Gabor analysis and algorithms}, pages 123--170. Birkh\"auser
  Boston, Boston, MA, 1998.


  
\bibitem{folland89}
G.~B. Folland.
\newblock {\em Harmonic Analysis in Phase Space}.
\newblock Princeton Univ. Press, Princeton, NJ, 1989.


\bibitem{G1} Gr\"{o}chenig, K.
  \newblock Multivariate Gabor frames and sampling of entire functions
  of several variables.
  \newblock {\em Appl. Comput. Harmon. Anal.} 31(2):218--227, 2011. 


\bibitem{G2}
K.~Gr{\"o}chenig.
\newblock {\em Foundations of time-frequency analysis}.
\newblock Birkh\"auser Boston Inc., Boston, MA, 2001.



\bibitem{G3} K.~Gr\"{o}chenig.
  \newblock Gabor frames without inequalities.
  \newblock {\em Int. Math. Res. Not.}  IMRN 2007, no. 23, Art. ID rnm111, 21 pp.


\bibitem{G4} K.~Gr\"{o}chenig.
  \newblock The mystery of Gabor frames.
  \newblock {\em J. Fourier Anal. Appl.} 20(4):865- 895, 2014.
  
% \bibitem{GL}  Gr\"{o}chenig,K. Lyubarskii, Yu. "Gabor (super)frames with Hermite functions", Math. Ann. 345 (2009) 267�286.

\bibitem{IR}
K.~Ireland and M.~Rosen.
\newblock{\em A classical introduction to modern number theory}, 
 volume 84 of {\em  Graduate Texts in Mathematics.} 
 \newblock  Springer-Verlag, New York, 1990.  

\bibitem{janssen95}
A.~J. E.~M. Janssen.
\newblock Duality and biorthogonality for {W}eyl-{H}eisenberg frames.
\newblock {\em J. Fourier Anal. Appl.}, 1(4):403--436, 1995.

\bibitem{GL86}
P.~Lelong and L.~Gruman.
\newblock {\em Entire functions of several complex variables}, volume 282 of
  {\em Grundlehren der Mathematischen Wissenschaften}.
\newblock Springer-Verlag, Berlin, 1986.

\bibitem{H1}  Hayman, W.
  \newblock The local growth of the power series: A survey of the
  Wiman-Valiron method.
  \newblock {\em Canad. Math. Bull.}  17:317--358, 1974. 

\bibitem{He} Heil, C.
  \newblock History and evolution of the density theorem
  for Gabor frames.
  \newblock {\em J. Fourier Anal. Appl.} 13(2):  113--166, 2007. 


%   \bibitem{L1}
% H.~J. Landau.
% \newblock Necessary density conditions for sampling and interpolation of
%   certain entire functions.
% \newblock {\em Acta Math.}, 117:37--52, 1967.


\bibitem{LLL}
A.~K. Lenstra, H.~W. Lenstra, Jr., and L.~Lov{\'a}sz.
\newblock Factoring polynomials with rational coefficients.
\newblock {\em Math. Ann.}, 261(4):515--534, 1982.

\bibitem{L1}
Y.~I. Lyubarskii.
\newblock Frames in the {B}argmann space of entire functions.
\newblock In {\em Entire and subharmonic functions}, pages 167--180. Amer.
  Math. Soc., Providence, RI, 1992.

\bibitem{LR00}
Y.~I. Lyubarskii and A.~Rashkovskii.
\newblock Complete interpolating sequences for {F}ourier transforms supported
  by convex symmetric polygons.
\newblock {\em Ark. Mat.}, 38(1):139--170, 2000.

\bibitem{MP00}
X.~Massaneda and P.~J. Thomas.
\newblock Interpolating sequences for {B}argmann-{F}ock spaces in {$\Bbb C^n$}.
\newblock {\em Indag. Math. (N.S.)}, 11(1):115--127, 2000.

\bibitem{OS}
J.~Ortega-Cerd\'a and K.~Seip.
\newblock  Beurling-type density theorems for weighted $L^p$ spaces of
entire functions. 
\newblock {\em J. Anal. Math. }
75 (1998), 247–266.

\bibitem{pap87}
D.~E. Papush.
\newblock The growth of entire functions with ``plane'' zeros.
\newblock {\em Teor. Funktsi\u\i\ Funktsional. Anal. i Prilozhen.},
  (48):117--125, 1987.

\bibitem{pap92}
D.~E. Papush.
\newblock Interpolation with discrete sets in {${\bf C}^l$}.
\newblock {\em J. Soviet Math.}, 59(1):666--674, 1992.
\newblock Translation of Teor. Funktsi{\u\i} Funktsional. Anal. i Prilozhen.
  No. 55 (1991), 113--124.

\bibitem{perelomov71}
A.~M. Perelomov.
\newblock Remark on the completeness of the coherent state system.
\newblock {\em Teoret. Mat. Fiz.}, 6(2):213--224, 1971.

\bibitem{PR13}
G.~E. Pfander and P.~Rashkov.
\newblock Remarks on multivariate {G}aussian {G}abor frames.
\newblock {\em Monatsh. Math.}, 172(2):179--187, 2013.

\bibitem{ron-shen97}
A.~Ron and Z.~Shen.
\newblock Weyl--{H}eisenberg frames and {R}iesz bases in ${L}\sb 2({\bR } \sp
  d)$.
\newblock {\em Duke Math. J.}, 89(2):237--282, 1997.

\bibitem{ronkin}
L.~I. Ronkin.
\newblock {\em Introduction to the theory of entire functions of several
  variables}.
\newblock American Mathematical Society, Providence, R.I., 1974.
\newblock Translated from the Russian by Israel Program for Scientific
  Translations, Translations of Mathematical Monographs, Vol. 44.

  \bibitem{S1}
K.~Seip.
\newblock Density theorems for sampling and interpolation in the
  {B}argmann-{F}ock space. {I}.
\newblock {\em J. Reine Angew. Math.}, 429:91--106, 1992.

\bibitem{S2}
K.~Seip and R.~Wallst{\'e}n.
\newblock Density theorems for sampling and interpolation in the
  {B}argmann-{F}ock space. {I}{I}.
\newblock {\em J. Reine Angew. Math.}, 429:107--113, 1992.

\bibitem{Simon2}
B.~Simon.
\newblock {\em Basic complex analysis}.
\newblock A Comprehensive Course in Analysis, Part 2A. American Mathematical
  Society, Providence, RI, 2015.

\bibitem{SB03}
T.~Strohmer and S.~Beaver.
\newblock Optimal ofdm design for time-frequency dispersive channels.
\newblock {\em Communications, IEEE Transactions on}, 51(7):1111--1122, July
  2003.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath,amsfonts,amsthm,amsopn,cite,mathrsfs}
\usepackage{epsfig,verbatim}
\usepackage{subfigure}
\newcommand{\gaul}{(\bZ +i\bZ )^d}
\newcommand{\duall}{\Lambda ^{\circ}}

\newcommand{\eps}{\epsilon}
\newcommand{\ccG}{\mathcal G}
\newcommand{\BF}{Bargmann-Fock space}
\newcommand{\bfd}{\cF _d}
\newcommand{\bC}{\mathbb C}
\newcommand{\bR}{\mathbb R}
\newcommand{\bZ}{\mathbb Z}
\newcommand{\bN}{\mathbb N}

\newcommand{\cd}{{\mathbb C}^d}
\def\intrd{\int_{\rd}}

\def\intrd{\int_{\rd}}

\newcommand{\cF}{\mathcal F}
\newcommand{\cE}{\mathcal E}
\newcommand{\cD}{\mathcal D}
\newcommand{\cZ}{\mathcal Z}
\newcommand{\cG}{\mathcal G}

\newcommand{\rd}{\bR^d}
\newcommand{\rdd}{\bR^{2d}}
\newcommand{\inv}{^{-1}}
\newcommand{\rems}{Rmarks}
\newcommand{\bba}{\mathbf a}
\newcommand{\bbb}{\mathbf b}
\newcommand{\mbZ}{\,\mathbb Z [i]}
\newcommand{\fif}{if and only if}
\def\lrd{L^2(\rd)}

\newcommand{\vs}{\bigskip}

\newtheorem{tm}{Theorem}[section]    
\newtheorem{lemma}[tm]{Lemma}
\newtheorem{prop}[tm]{Proposition}
\newtheorem{cor}[tm]{Corollary}

\newcommand{\tfs}{time-frequency shift}

\newcommand{\beq}{\begin{equation}}
\newcommand{\eeq}{\end{equation}}

\begin{document}
\begin{abstract}
We give a general construction of entire functions in $d$ complex
variables that vanish on a lattice of the form $\Lambda = A (\bZ +
i\bZ )^d$ for an invertible complex-valued matrix. As an application
we exhibit a class of lattices of density $>1$ that fail to be a sampling set for the
Bargmann-Fock space in $\bC ^2$. By using an equivalent real-variable
formulation, we show that these lattices fail to generate a Gabor
frame. 
\end{abstract}

\title[Lattice Sampling of Entire Functions]{
Sampling of Entire Functions of Several Complex Variables on a Lattice and 
  Multivariate Gabor Frames }
\author{Karlheinz Gr\"ochenig}
\address{Faculty of Mathematics \\
University of Vienna \\
Oskar-Morgenstern-Platz 1\\
A-1090 Vienna, Austria}
\email{karlheinz.groechenig@univie.ac.at}
 \author{Yurii Lyubarskii}
\address{Department of Mathematical Sciences, Norwegian University of
 Science and Technology, NO-7491, Trondheim, Norway} 
\email{yura@math.ntnu.no} 
\subjclass[2010]{42C15, 33C90, 32A30, 94A12}
\date{}
\keywords{Gabor frame, Gauss function, lattice, Weierstrass
  sigma-function, interpolating function,  entire functions of several variables}
\thanks{K.\ G.\ was
  supported in part by the  project 31887-N32 the
Austrian Science Fund (FWF)}
\maketitle

\section{Introduction}

We study the sampling problem in the Bargmann-Fock space of several
complex variables and the related construction of Gabor frames with a
Gaussian window. Our main point is the restriction to sampling on a
lattice in $\cd$ and the consequences resulting from  the additional
invariance properties. 

The first problem is a sampling problem for entire functions in
several complex variables.  Recall that 
the Bargmann-Fock space $\cF ^2 _d$ consists of all 
entire functions of $d$ complex variables $z= (z_1, \dots, z_d) \in
\cd $ with finite norm  
\begin{equation}
  \label{eq:n1}
\| F\|_{\cF^2_d}^2 =  \int _{\cd } |F(z)|^2 e^{-\pi |z|^2} \, dz . 
\end{equation}
A set  $\Lambda \subseteq \cd $ is called  a {\em sampling set} for
$\cF ^2_d$, if, for some constants $A,B>0$, 
\beq
\label{eq:n2}
A \|F\|_{\cF ^2_d}^2 \leq \sum_{\lambda \in \Lambda} |F(\lambda)|^2 e^{-\pi |\lambda|^2}
 \leq B \|F\|_{\cF _d^2}^2, \quad \forall F\in \cF ^2_d. 
\eeq

Our second question deals with the spanning properties of \tfs s of
the 
 Gaussian function $\phi(x)=
 \exp(-\pi |x|^2)$, 
 $x\in\rd$, and thus is a  problem for functions on $\rd $.  Let   $\lambda=\xi+i \eta \in \cd$ with  $\xi,\eta \in \rd$   and 
$$
(\pi_\lambda \phi)(x)=
e^{2i\pi <\eta, x>}e^{- \pi |x-\xi|^2}
$$
be the  corresponding \tfs\ of $\phi $ by $\lambda $. 
Given a discrete set $\Lambda \subseteq \cd $, we denote the set of
\tfs s along $\Lambda $ by 
$$
\ccG(\Lambda)= \{\pi_\lambda \phi : \lambda\in \Lambda \} \, ,
$$
which is usually  called a  Gabor family.   We then say that  $\ccG(\Lambda)$ is {\em a frame} for   
$L^2(\rd)$,  if  for some constants $A  , B>0$
\beq
A \|f\|^2_{L^2(\rd)}  \leq \sum_{\lambda\in \Lambda} |\langle f, \pi_\lambda \phi \rangle |^2
\leq B  \|f\|^2_{L^2(\rd)}, \quad \forall  f \in  L^2(\rd).
\eeq
We are interested in the frame property of $\ccG(\Lambda)$ in
$L^2(\rd)$. This is a ``real-variable'' problem about functions in
$\lrd $. 

It is well-known that the Bargmann transform maps the \tfs s $\pi
_\lambda \phi$ to the normalized reproducing kernel of $\cF
^2_d$~\cite{bargmann61}. Therefore these two problems are equivalent
via the Bargmann transform:
 
 {\em The system $\cG(\Lambda)$  forms a frame in $L^2(\bR^d)$,  if
   and only if  $\Lambda$ is sampling in $\cF^2_d$.}
  See Section~\ref{duality1} for a detailed description.

 In  the case of one variable  $d=1$ the sampling property in $\cF
 ^2_1$ and  frame 
property of $\cG (\Lambda )$ can be completely characterized  in  terms of the 
density of $\Lambda $  by the results in \cite{L1, S1,
  S2}. Precisely, a separated
set $\Lambda \subseteq \bC $ is a sampling set for $\cF ^2_1$, \fif\
its  (lower) Beurling $D^-(\Lambda )$ is greater than $ 1$. See  also
the discussion in \cite{G2,G4,He}.
In the multivariate case  the density condition $D^-(\Lambda ) >1$  is 
necessary (see \cite{OS,MP00}), but is far from sufficient. Sufficient conditions in terms of
a covering density are given in~\cite{fg89jfa,MP00}, but they imply a
large  Beurling density. 

For more detailed results,  additional arithmetic conditions are
required. 
For this reason we restrict our attention exclusively to  lattices. 
As in \cite{G1} we consider only complex lattices.  By a complex
lattice we understand 
a   lattice of the form
$$\Lambda = A  (\bZ + i \bZ )^d = A \mbZ ^d
$$
for some $A\in \mathrm{GL}\, (d,\bC )$. Throughout  we will write $\mbZ = \bZ +
i\bZ $ for the ring of Gaussian integers in $\bC$. 
Since $\Lambda $ is a discrete subgroup of $\cd $, the corresponding
sampling set and the Gabor family $\cG (\Lambda )$ possess 
 an additional   structure.

To answer the   question about  Gabor frames over a lattice,
one applies the  fundamental
duality theory~\cite{feichtinger-kozek98,janssen95,ron-shen97} (see also
Section~\ref{duality1} below).  One of the key points is relation between the sampling problem in the space 
$\cF^2_d$ and the uniqueness problem in the space $\cF^\infty_d$,
which  consists of entire functions $F$ on 
$ \cd$ such that 
$$
\|F\|_{\cF^\infty_d}= \sup_{z \in \cd} |F(z)| \, e^{-\frac {\pi |z|^2}2} < \infty.
$$
After a suitable reformulation, we see that the
construction of Gabor frames is  intimately connected to two fundamental problems about
entire functions. 
 

 
 \begin{itemize}
 \item[(i)] Construct  an entire  function $\sigma _\Lambda  $ with possibly smallest growth that vanishes on
   $\Lambda $, which in analogy to the one-dimensional case we call a
   sigma-type function for $\Lambda $.
 \item[(ii)] Construct an entire function $\tau _\Lambda $ that is interpolating
   on $\Lambda $, i.e., $\tau _\Lambda (\lambda ) = \delta _{\lambda
     ,0}$ for $\lambda \in \Lambda $.  
 \end{itemize}

 In dimension $d=1$ the above problems on $\mbZ $  are solved by the
 classical Weierstrass  $\sigma$-function  
$$
\sigma (z)=z\prod_{m,n\in \bZ, \  (m,n)\neq (0,0)}\left ( 1-\frac z {m+in} \right )
e^{\frac z {m+in} +\frac 1 2\frac  {z^2}{(m+in)^2}} \, , \quad \,\, \tau(z)=\frac {\sigma(z)}z \, .
$$
Clearly, $\sigma $  is an entire function   with $\mbZ$
as its  zero set, and $\tau $ in interpolating on $\mbZ $. 

By contrast, in  several complex variables the above
questions are rather unusual.   The zero set of an
entire function of $d>1$ variables is always an analytic
manifold. Although it is possible to extend the construction of
the Weierstrass product  to obtain entire functions whose zero set is
a given analytic hypersurface ~\cite{GL86,ronkin}, this
construction sheds no light 
on the search of  sigma-type functions on a lattice. Only few results about
interpolation with discrete sets are known, see, e.g.,~\cite{pap87,pap92}.

Our first contribution is a general recipe for the  construction of   sigma-type functions and interpolating functions
associated to an arbitrary complex lattice. The idea  builds essentially on the
one-dimensional machinery and  yields  a  special class of sigma-type  functions, whose
zero set  is a union of analytic planes that contain the original
lattice.   Entire
functions on $\bC ^2$ with plane zeros also  play an important role in
~\cite{LR00}. In principle this idea works for arbitrary dimensions, but
we will restrict ourselves to entire functions of \emph{two}
variables.

Our second contribution is the application of the general construction
of sigma-type functions 
to show that certain complex lattices 
fail to yield sampling sets for $\cF ^2_2$.  This requires the
control  of    the growth of the
sigma-type function so that it is in $ \cF ^2_d$.  For the
application to sampling in Bargmann-Fock space the goal is therefore
to find sigma-type functions with small growth. 

To provide an idea of the main  construction, we consider a model example 
that inspired our general construction.

Let 
\begin{equation}
 \label{matrixaex}
A=\begin{pmatrix}
  1 & 1/2 \\
0 & \sqrt{3}/2
\end{pmatrix},  \quad   \Lambda=A   \mbZ^2.
\end{equation}
This lattice is the complexification of the usual hexagonal lattice in
$\bR ^2$. 
Our main construction then suggests the following sigma-type function
for $\Lambda $: 
\begin{equation}
  \label{eq:n1}
\sigma _\Lambda (z_1,z_2)= \sigma(z_1) \, \sigma (\frac{z_2}{\sqrt 3} -
\frac{1}{2} ) \, \sigma (\frac{z_2}{\sqrt 3} -
\frac{i}{2}  ) \,  \sigma (\frac{z_2}{\sqrt 3} -
\frac{1+i}{2}  ) e^{2\pi (1-i) z_2} \, .
\end{equation}

It is easy to check that    $\sigma _\Lambda $ vanishes on $\Lambda $ and satisfies the growth estimate
  $|\sigma _\Lambda (z)| \leq C e^{\pi |z|^2/2}$.

The general characterization of lattice sampling sets  
(Proposition~\ref{chargauss}) then  implies the following result: 
\begin{center}
\emph{$\Lambda $ fails to be a sampling set for $\cF ^2_2$.}  
\end{center}



By contrast,  the  sigma-type function
$$
\sigma _2  (A\inv z) = \sigma  (z_1-\frac{z_2}{\sqrt{3}}) \sigma (\frac{2z_2}{\sqrt{3}})
$$
also vanishes on $\Lambda $, but it grows much too fast to be of use in the analysis of $\cF ^2_2$.

A small modification of \eqref{eq:n1} yields the interpolating function
\begin{equation}
  \label{eq:n2}
\tau _\Lambda  (z_1,z_2)= \frac{\sigma(z_1)}{z_1} \, \sigma (\frac{z_2}{\sqrt 3} -
\frac{1}{6} ) \, \sigma (\frac{z_2}{\sqrt 3} -
\frac{i}{6}  ) \sigma)(\frac{z_2}{\sqrt 3} -
\frac{1+i}{6}  ) e^{\pi (1-i)/\sqrt{3} z_2} \, .
  \end{equation}
 Although $\tau _\Lambda  \not \in \cF _2^2$, it satisfies the growth
 rate $\phi _\Lambda (z) \leq C e ^{\pi |z|^2/2}$. This suffices to
 derive a weak   Lagrange interpolation formula. 
See Theorem~\ref{weaksamp}. 

This example is rather puzzling. In dimension $d=1$ the hexagonal
lattice $A\bZ ^2$  has density $2/\sqrt{3}>1$ and generates a Gabor
frame $\cG (A\bZ ^2)$ with certain  optimal
features~\cite{SB03,FS17}.
By contrast in dimension $d=2$, $A\mbZ ^2$  fails to generate a Gabor
frame and a sampling set for $\cF ^2_2$, although it has density $4/3>1$.

Currently the investigation of sampling the Bargmann-Fock space on
lattices is poorly understood, and still amounts to the investigation
of examples and counter-examples. We hope that this article will stir
some interest among  the experts in several complex variables and that it
will inspire a deeper analysis of the problem. 

 \medskip

The article is organized as follows.  In Section~2 we collect the
background material about the connection between sampling in
Bargmann-Fock space and Gabor frames, the basic information about the
Weierstrass sigma-function, and a normalized representation of
lattices by means of Minkowski reduced bases. Section~3 contains the
main construction of sigma-type functions and interpolating functions
for complex lattices. In Section~4 we use this class of sigma-type
functions to show that certain ``natural'' lattices of density $>1$
fail to be sampling in Bargmann-Fock space. In Section~5 we prove a
weak Lagrange interpolation formula for certain lattices of density
$\geq 1$.


\section{Sampling, Sigma Functions, and Lattices}

\subsection{Sampling in Bargmann-Fock space and Gabor frames }     
\label{duality1}
The lattice structure  leads to special
criteria  for a set $\Lambda $ to be sampling. We emphasize that these are unavailable
for arbitrary sets of points. 


 

The  relation between the sampling property of the lattice $\Lambda$
and the frame property of the system $\ccG(\Lambda)$ 
 is  summarized in the following statement. For its formulation we
 need the  adjoint lattice $\duall =(A^*)\inv
\mbZ^d $  of $\Lambda = A \mbZ ^d$ 

\begin{prop} \label{chargauss}
 
For a  lattice $\Lambda = A \mbZ ^d  \subseteq \bC ^d$  the following are
equivalent: 

(i) $\ccG(\Lambda)$ is a frame in $L^2(\rd)$;

(ii) $\Lambda $ is a  sampling set for $\cF^2_d$; 

(iii)  There exists an interpolating  function $G \in \cF^2_d $ for
$\Lambda ^\circ$  satisfying  the Bessel property, 
i.e.,  
\begin{equation}
  \label{eq:22}
  G(\mu ) = \delta _{\mu ,0} , \qquad \text{ for all } \mu \in \Lambda
  ^\circ \, , 
\end{equation}
and  $F \to \big( \langle F , e^{\pi \bar{\lambda } \cdot z} G(z − λ)
\rangle e^{-\pi |\lambda |^2/2}\big)_{\lambda \in \Lambda } $ maps
$\cF ^2_d$ to
$\ell ^2(\Lambda )$.

(iv) $\Lambda $ is a set of uniqueness for $\cF ^\infty_d $, i.e., if $F
\in \cF ^\infty_d $ and $F (\lambda ) = 0$ for all $\lambda \in \Lambda
$, then $F \equiv 0$. 
\end{prop}

Proposition~\ref{chargauss} explains the fundamental importance  of  sigma-type
functions and interpolating functions in $\cF ^2_d$  for the theory of
Gabor frames. 
It follows from this proposition that a function $F\in \cF ^2_d$ can 
be recovered from its samples $\{F(\lambda)\}_{\lambda\in \Lambda}$
by mean of the interpolating function $G$  constructed   for $\duall $, not for $\Lambda
$!  Namely \footnote{We always  use the "real" inner product: $z\cdot
  w=\sum _{j=1}^d z_j w_j$ for
$z=(z_j), \ w=(w_j) \in \bC ^d$. }
\begin{equation}
  \label{eq:n5}
  F(z) = \sum _{\lambda \in \Lambda } F(\lambda ) e^{\pi \bar{\lambda
    }\cdot z} G(z-\lambda ) e^{-\pi |\lambda |^2/2} \, .
\end{equation}


The connection (i) and (ii) is classical. Let $f\in \lrd $ and 
\begin{equation}
  \label{eq:n9}
  Bf(z) = 2^{d/4} \, e^{-\pi z\cdot z /2} \, \intrd f(x) e^{-\pi |x|^2 + 2\pi x\cdot z} \, dx 
 \end{equation}
be the Bargmann transform of $f$. Then $B$ is unitary from $\lrd $
onto the Bargmann-Fock space $\cF ^2_d$, e.g., by
\cite{folland89}. Moreover, 
 the Bargmann transform maps
the \tfs\ $2^{d/4}\, \pi _\lambda \phi $ to the normalized  reproducing kernel $e^{\pi
  \bar{\lambda } \cdot z} e^{-\pi |\lambda |^2/2}$ of the Fock
space. Therefore 
\begin{equation}
  \label{eq:n11}
  \langle f, \pi _\lambda \phi \rangle _{L^2} =
\langle Bf, B(\pi _\lambda \phi )\rangle _{\cF ^2_d} = Bf(\lambda )
e^{-\pi |\lambda |^2/2} \, .
\end{equation}
See ~\cite{folland89,G2} and the original
literature~\cite{bargmann61}.

The equivalence of (ii) and (iii) seems beyond the realm of complex
analysis and is due to the invariance properties of lattices. It shows
that the problem of sampling on a given lattice $\Lambda $ is
equivalent to an interpolation problem of the adjoint lattice $\Lambda
^\circ $. This statement  is  part of the duality theory of
Gabor frames~\cite{feichtinger-kozek98,janssen95,ron-shen97}. 

The
equivalence of (ii) and (iv)  follows from  one of  the characterizations of
Gabor frames without 
inequalitites~\cite{G3} via the Bargmann transform. 
\medskip

\subsection{Sigma-type functions for lattices}


 Our  main  tool for the construction of sigma-type functions in $\cd
 $  is  the classical Weierstrass  $\sigma$-function  of {\em one} variable 
$$
\sigma (z)=z\prod_{\lambda \in \mbZ }\left ( 1-\frac z {\lambda }
\right ) e^{\frac z {\lambda } +\frac 1 2\frac  {z^2}{\lambda ^2}}.
$$
We refer to \cite{A1,Simon2} for basic  properties of $\sigma(z)$. In
particular, $\sigma$  is an entire function   with $\mbZ$ as the zero set.
 
In addition , see e.g. \cite{H1}, for each $\eps>0$
\begin{equation}
\label{hayman}
| \sigma(z)| \asymp e^{\pi |z|^2/2}, \quad \mathrm{dist}(z,\mbZ)>\eps.
\end{equation}
We note that $\sigma \in \cF ^\infty _1 \setminus \cF ^2_1$. By
Proposition~\ref{chargauss} the lattice of Gaussian integers $\mbZ $
fails to be a sampling set for $\mbZ $. This fact was already proved
in~\cite{bargmann71,perelomov71}

For $z=(z_1, \ldots \ , z_d)\in \cd$ we set
\begin{equation}
  \label{eq:c1}
  \sigma _0 (z) = \prod _{j=1}^d \sigma (z_j) \, .
\end{equation}
Then the function 
\begin{equation}
  \label{eq:c2}
  \sigma _A (z) = \sigma _0(A\inv z)
\end{equation}
vanishes on $\Lambda $ and satisfies the growth estimate 
 \begin{equation}
 \label{eq:c2a}
  |\sigma _A  (z) | \leq C e^{\pi \|A\inv \|_{op}^2 |z|^2/2}\, ,
\end{equation}
where as usual $\|A\|_{op}$ denotes the largest singular value of $A$.

Similarly we construct the {\em interpolating } function for $\Lambda=
A \mathbf{Z}^d$. 
 We define $\tau(z)=z^{-1}\sigma(z)$ for $z\in \bC$ and
 \begin{equation}
 \label{eq:c2c}
 \tau_0(z)=\prod_{j=1}^d \tau (z_j); \ \tau_\Lambda(z)=\tau_0(A^{-1}z) \ \mbox{for}  \  z=(z_1,\ldots \ , z_d)\in \bC^d. 
 \end{equation}
For example, consider  the lattice $\alpha \mbZ  \subseteq \bC $ with
adjoint lattice $\duall = \alpha \inv \mbZ $. Then $\duall $
possesses the interpolating function $\tau (\alpha z)$, which belongs
to $\cF ^2_1$, \fif\ $0<|\alpha | <1$. By Proposition~\ref{chargauss}
$\alpha \mbZ  $ is a sampling set for $\cF ^2_1 $, \fif\ $|\alpha | <1$. 
This is a baby version of  the complete characterization of
one-dimensional sampling sets in~\cite{L1, S1,S2}.  
\medskip


\subsection{Lattice reduction and Minkowski-reduced bases}

Clearly, the functions $\sigma _\Lambda $ and $\tau _\Lambda $ depend
on the generating matrix $A$, or equivalently, on the choice of a
basis of $\Lambda $. Our first task is to choose a suitable basis of
$\Lambda $, so that $\|A\inv \|_{\mathrm{op}}$ is small. 
One of possible recipes is to choose
a basis of $\Lambda $ which consists of the shortest possible  vectors. 
 
Precisely, choose  vectors $\mathbf{a}_j\in \Lambda$,
such that $\|\bba _1 \| = \min \{\|\lambda \| : \lambda \in \Lambda
\}$ and 
$$
\|\bba _j \| = \min \{ \|\lambda \| : \lambda \in \Lambda , \lambda
\not \in \mathrm{span}\, [\bba _1, \dots \bba _{ j-1}] \} \, .
$$
Such a  basis is called 
Minkowski reduced for $\Lambda $ ~\cite{LLL}, and  satisfies
$$
\|\bba _1\| \leq \| \bba _2 \| \leq \dots \leq  \| \bba _d \| \, . 
$$
After setting $A = \big(  \mathbf{a} _1 \,\,  \mathbf{a} _2 \,\, \dots
\,\, \mathbf{a} _d \big)$,  
the lattice is $\Lambda = A \mbZ ^d$. 

 We may  write $A=US$, where $U$ is a
 unitary matrix   and $S$ is an upper triangular matrix with real
 values on the diagonal and 
 with columns $\bbb _j$ (by  QR-factorization or  Gram-Schmidt
 orthogonalization). Since the Fock space $\cF ^2_d$ is invariant
 under a unitary transformation  $U$  of coordinates, a set $\Lambda $
 is sampling, \fif\ $U\inv \Lambda $ is sampling. Replacing $\Lambda =
 A\mbZ ^d$ by $U\inv \Lambda = S \mbZ ^d$, we may therefore
 assume without loss of generality that  $\Lambda = S \mbZ ^d$ for an
 upper triangular matrix $S$. 
 
Since $U$ preserves lengths, the columns
of $S$ are also Minkowski-reduced basis, and  $\|\bbb _1\| \leq \| \bbb _2 \|
\leq \dots \leq \| \bbb _d \| $.  The upper triangular matrix $S$ is of
the form 
\begin{equation}
  \label{eq:r1}
S = (s_{jk})= 
\begin{pmatrix}
   \gamma _1 & \ast & \dots & \ast    \\
0 & \gamma  _2 &   \dots & \ast  \\
\vdots \\
 0 &  \dots  & 0 & \gamma _d 
\end{pmatrix} \, ,
\end{equation}
and its entries satisfy the additional conditions 
\begin{align}
s_{j,k}  &= 0 \qquad k<j, \quad s_{kk}  = \gamma _k
  >0,       \label{eq:n3} \\
 |\mathrm{Re} & \, s_{j,k} |, \quad |\mathrm{Im}\, s_{j,k} |   \leq
  \frac{s_{j,j}}{2} \, . \end{align}
 In   \cite{G1} we proved the following result for general
 upper-diagonal matrices in arbitrary dimension.  

\begin{prop}
  \label{known}
  Let $\Lambda = S\mbZ ^d $ with an upper triangular matrix $S$. 

(i)   If $\gamma _j <1$ for $j=1, \dots ,d$, then $\Lambda $ is a sampling
  set for $\cF ^2_d$ (and $\cG (\Lambda )$ is a frame for $\lrd $).

 (ii) If $\gamma_d\geq 1$, then $\Lambda $ is not sampling and
 $\ccG(\Lambda)$ is not  a frame.  
\end{prop}

The proof goes roughly as follows. The adjoint lattice is  $\Lambda ^\circ
= (S^*)\inv \mbZ ^d$, where $(S^*)\inv $ is a lower trigonal matrix
with diagonal $ (\lambda _j \inv )_{j=1, \dots ,d}$. 
Then  the function  
$$
F_0(z_1,z_2, \dots , z_d)= \prod _{j=1}^d \frac {\sigma(\gamma_j z_j)}{z_j}
$$
belongs to $\cF^2_d$ and solves the interpolation problem $(iii)$ in Proposition \ref{chargauss}.
If $\gamma _d\geq 1$, then the  function $F_1(z_1,z_2, \dots , z_d):=
\sigma(\gamma_d^{-1} z_d)$ belongs to $\cF^\infty _d$ and vanishes on
$\Lambda$. By Proposition \ref{chargauss} $\Lambda $ fails to be a
sampling set. 

A different class of examples is discussed in ~\cite{PR13}.

From now on we deal with explicit constructions in two complex
variables.  We may assume without loss of
generality that $\Lambda $ is determined by the matrix
\begin{equation}
\label{matrixa}
A =
\begin{pmatrix}
  \gamma _1 & \beta \\
0 & \gamma _2
\end{pmatrix}
\end{equation}
with 
\begin{equation}
\label{eq:d9a}
\gamma _1, \gamma _2 >0, \quad  \gamma _1 ^2 \leq |\beta |^2 + \gamma _2^2. 
\end{equation}
Since the basis is reduced, we  have 
\begin{equation}\label{eq:d9b}
  |\mathrm{Re}\, \beta|, \quad |\mathrm{Im}\, \beta|   \leq \frac{\gamma
  _1}{2} .
\end{equation}


\section{A  Construction of Sigma-Type Functions and Interpolating
  Functions via Sublattices } 

To go beyond Proposition~\ref{known}, we need a more involved
recipe for sigma-type and interpolating functions. 
In  this section we provide such  a  general construction  
inspired  by the model 
example \eqref{eq:n1}. The goal is to produce sigma-type  functions or
interpolation functions   with   small   
growth. 

Given a lattice   $\Lambda = A \mbZ^2 $  in $\bC^2 $, we will use the
following master plan to construct a   sigma-type  function 
 function or an interpolating function  on $\Lambda $ via a sublattice. 
\begin{enumerate}
\item Construct a sublattice $\Gamma \subseteq \Lambda  $ that
  possesses an orthogonal basis (or possibly a nearly orthogonal
  basis). 
\item Construct a sigma function $\sigma _\Gamma $  and an
  interpolating function $\tau _\Gamma $ on
  $\Gamma $ according to~\eqref{eq:c2} and \eqref{eq:c2c}.
\item Determine a suitable set  of coset representatives $ \Lambda  /
  \Gamma $. 
\item The sigma function  and the interpolating  function on $\Lambda
  $    will be a suitable product of shifts
  of each of the factors of $\sigma_\Gamma $ and $\tau _\Gamma
  $.  
\end{enumerate}

\subsection{Sublattices and their cosets}
Following the  outline above, we first describe sublattices and their
cosets in $\bC ^2$. 

 Every (complex)  sublattice of $\mbZ^2 $ is of
the form $B \mbZ^2 $ for an invertible  matrix $2\times 2$-matrix
$B$ with entries in $\mbZ$ (in short, $B\in
\mathrm{GL} (2, \mbZ )$).   
Consequently, every sublattice of a lattice $\Lambda = A \mbZ^2
\subseteq \bC ^2 $ is
of the form 
$$
\Gamma = A B \mbZ^2
$$
 for some $B\in \mathrm{GL} (2, \mbZ )$.

Clearly, we can represent the full lattice $\Lambda $ as a union of
shifts of the sublattice $\Gamma $ as follows $\Lambda = \bigcup
_{j=1}^n (\delta _j + \Gamma )$ for some lattice points $\delta _j\in
\Lambda $. In fact, the shifts $\delta _j$  are the representatives
 of the quotient $\Lambda / \Gamma $. 
We will use the following explicit  parametrization of $\mbZ ^2/ B
\mbZ ^2$ and hence of  $\Lambda  / \Gamma $.             

\begin{lemma}\label{l3}
  Let  $B=\big(
  \begin{smallmatrix}
    a & c \\ b &d
  \end{smallmatrix}
\big)$ with   entries $a,b,c,d \in \mbZ $. Let $\gamma = \mathrm{gcd}
(a,c)\in \mbZ$~\footnote{Note that over $\mbZ $ the greatest common divisor is
only   determined up to multiplication with $\pm 1, \pm i$. We refer the reader to \cite{IR}
for the facts on division in $\mbZ$.} and  $\Delta
= \det B  = ad-bc  $. 
Let $Q=  \big([0,1) + i[0,1)\big)\subseteq \bC $.  Then  the set
$$
\cD = \{ (\delta _1,\delta _2 )\in \mbZ^2: \delta _1\in \gamma Q \cap
\mbZ , \delta _2 \in \tfrac{\Delta}{\gamma} Q \cap \mbZ  \} 
$$ 
is a set of coset representatives for $\mbZ^2 / B\mbZ ^2$. 

In particular, if $B$ possesses \emph{real-valued entries} $a,b,c,d
\in \bZ $,  and $a$ and $c$ are relatively prime over $\bZ $, then 
$$
\cD = \{ (0,\delta );   \delta= \alpha+i \alpha' ), \  0\leq \alpha, \alpha ' < |\det B|\} 
$$
is a set of coset representatives of $\mbZ^2 / B\mbZ^2 $.
\end{lemma}

\begin{proof}
We observe that the set $Q$ is a (half-open)  square in $\bC $ and
that $(Q-Q) \cap \mbZ  = \{0\}$, where  $Q-Q = (-1,1) + i (-1,1)$ is
the difference set. 
We write $a= \gamma a', c = \gamma c'$ for $a',c'\in \mbZ $, and note
that $\Delta = \gamma (a' d - b c') $, so that $\Delta /\gamma \in
\mbZ $. 

Assume that $ \delta _1, \delta _1'  \in \gamma Q \cap \mbZ$, $ \delta
_2, \delta _2'  \in \frac{\Delta}{\gamma }  Q \cap \mbZ$ and  that
$(\delta _1- \delta _1', \delta _2 - \delta _2' )^T = B
(k,l)^T$ for some $k,l \in \mbZ $, in other words $(\delta _1, \delta _1')$ and
$(\delta _2, \delta _2')$ represent the same coset.  Then 
$$
\begin{pmatrix}
  a& c \\ b & d 
\end{pmatrix}
\begin{pmatrix}
  k \\ l 
\end{pmatrix}
=
\begin{pmatrix}
  ak+cl \\
bk+dl
\end{pmatrix} = \begin{pmatrix}
\delta _1-\delta _1' \\
\delta _2-\delta _2' 
\end{pmatrix} \, .
$$
The first coordinate is   
$$
\delta _1-\delta _1' = ak+cl = \gamma (a'k+c'l) \in \gamma (Q-Q) \cap
\mbZ  \, , $$
and therefore the Gaussian integer $a'k+c'l $ is in $Q-Q$, which
implies that $a'k+c'l = 0$. 
Since $\mathrm{gcd} (a', c') = 1 $ (up to multiplication by $\pm 1,
\pm i$), it follows that $l=Na', k=-Nc'$ for some $N\in \mbZ $.  

Now the second coordinate of $B(k,l)^T$ is
$$
bk+dl = N(-bc' + a'd) = N \frac{\Delta}{\gamma} \in \frac{\Delta}{\gamma} (Q-Q) \cap \mbZ  \,
.
$$
Therefore 
$N\in (Q-Q) \cap (\mbZ
)$. This implies that $N=0$ and thus $k=l=0$. Altogether we have shown
that $(\delta _1, \delta _1') = (\delta _2, \delta _2')$. It is easily
verified that the
cardinality of  $\gamma Q \cap \mbZ $ is $|\gamma |^2$ for $\gamma \in
\mbZ $, therefore the cardinality of $\cD $ is $ |\det
B| ^2$. Consequently  $\cD $ is a complete
set of  representatives of $\mbZ ^2 / B \mbZ ^2$. 

\end{proof}




\subsection{A  construction of sigma-type  functions} \label{4.3}

We begin with the  calculation of $\sigma _\Gamma $ and $\tau _\Gamma
$ on the cosets $\nu +\Gamma $ of $\Gamma $ in $\Lambda $. 
Let 
$B=\big(
  \begin{smallmatrix}
    a & c \\ b &d
  \end{smallmatrix}
\big)$,  $ \Delta= \det B$ and $\Gamma = AB \mbZ ^2$ be the
sublattice of $\Lambda = A \mbZ ^2$. In view of the explicit examples,
we  assume furthermore that 
$|\mathrm{gcd }\, (a,c) | = 1$. 

Let, as before, $\sigma(z)$ be the classical Weierstrass
$\sigma$-function for $\mbZ$. For $z=(z_1,z_2)\in   \bC$
we denote $\sigma \otimes \sigma (z)=\sigma (z_1)\sigma(z_2)$ and set 
$$
\sigma _\Gamma  (z) = (\sigma \otimes \sigma ) \Big( (AB)\inv
z\Big).
$$

We first evaluate $\sigma _\Gamma $ on $\Lambda $. Since
$|\mathrm{gcd}(a,c)| = 1$, the representatives of $\mbZ ^2/ B\mbZ^2$
can be chosen to be $(0,\delta )$, and thus a  general lattice
point  $\lambda \in \Lambda = A\mbZ^2$ can be written as 
\begin{equation}
\label{eq:q1}
\lambda =  A \Big( B
\begin{pmatrix}
  k \\ l 
\end{pmatrix}
+
\begin{pmatrix}
  0 \\ \delta 
\end{pmatrix}
\Big) \, 
\end{equation}
 for some $ k,l \in
\mbZ $ and $(0,\delta ) \in \cD $. 
Consequently  
\begin{align} 
\sigma_\Gamma (\lambda ) &= (\sigma \otimes \sigma ) \Big( (AB)\inv
  \big(AB (k,l)^T + A (0,\delta)^T\big) \Big) \\
&= (\sigma \otimes \sigma ) \Big(
\begin{pmatrix}
  k \\ l 
\end{pmatrix}
+ B\inv
\begin{pmatrix}
  0 \\ \delta 
\end{pmatrix}
\Big) \notag \, .
\end{align}
 Since 
$$ B\inv \begin{pmatrix}
  0 \\ \delta 
\end{pmatrix}
= \frac{1}{\Delta}
\begin{pmatrix}
  d & -c \\ -b &a
\end{pmatrix}
\begin{pmatrix}
  0 \\ \delta 
\end{pmatrix}
=
\begin{pmatrix}
  -c\delta /\Delta \\ a\delta / \Delta 
\end{pmatrix} \, ,
$$
the sigma function of the sublattice $\Gamma $ evaluated on $\Lambda  $
is
\begin{equation}
  \label{eq:d3}
  \sigma_\Gamma (\lambda ) = \sigma \big(k- \frac{c\delta }{\Delta}\big)
  \sigma \big(l+ \frac{a\delta }{\Delta}\big)   \, .
\end{equation}

Here is our  key  observation:

\emph{If   $\frac{\Delta}{c}$  divides $\delta$, then $\frac{c\delta
}{\Delta}\in \mbZ$ and $\sigma (k- \frac{c\delta
}{\Delta}) = 0$, thus  $\sigma _\Gamma \big(
\big( \begin{smallmatrix}
  0 \\ \delta
\end{smallmatrix} \big)
+\gamma \big) = 0$ for all $\gamma \in \Gamma $}.

Similarly,  if
$\frac{\Delta}{a} $ divides $\delta $, then $  \sigma (l+
a\delta / \Delta )  =0$. 
Thus a single factor of $\sigma_\Gamma$ in the product~\eqref{eq:d3} vanishes on the whole coset
$(0,\delta) +\Gamma$ and we need not include additional factors   
to $\sigma_\Lambda$  in order to annihilate this coset.
We will therefore try to choose a
sublattice $\Gamma $  such
that $a$ or $c$ divides $\det B  = \Delta $ and  $\Delta / c$ is
small. 

To take care of the cosets that do not vanish in this way, we use
(one-dimensional) Fock shifts. For $z,\zeta \in \bC $ let
\begin{equation}
\label{eq:d9bca}
\beta _\zeta f(z) = e^{\pi \bar{\zeta}z-\pi |\zeta|^2/2} f(z-\zeta )
\, .
\end{equation}
 We have 
$$
e^{-\frac \pi 2 |z|^2} |\beta_\zeta f(z)|= e^{-\frac \pi 2 |z-\zeta|^2}|f(z-\zeta)|,
$$
so  $\beta
_\zeta $ is a unitary operator on $\cF ^2_1$ and an isometry
on $\cF ^\infty _1$.    In particular, 
together with \eqref{hayman} this yields
\begin{equation}
\label{eq:d9bbb}
|\beta_\zeta\sigma(z)| \asymp e^{\pi |z|^2/2}\, , \,\, \text{ if } \,\, \ \mbox{dist}(z, \mbZ+\zeta)>\eps.
\end{equation}





For $z= (z_1,z_2) \in \bC ^2$ let $p_1(z_1,z_2) = z_1$ and
$p_2(z_1,z_2) = z_2$ be the projections onto the first and second
coordinate of $z$. In this notation
$\sigma _\Gamma (z) = \sigma \big( p_1((AB)\inv z)\big) \, \sigma \big(
p_2((AB)\inv z)\big)$.

\vs
We split the construction of a sigma-type function for $\Lambda $  into several steps.

\textbf{Step 1.} Partition the coset representatives  $\cD $ into disjoint subsets
\begin{equation}
  \label{eq:vm1}
\cD =   \{ (0,\delta ) \in \cD : \frac{\Delta }{c} |
\delta \} \cup \cE _1 \cup \cE _2  = \cE _0 \cup \cE _1 \cup \cE _2 \,
,  
\end{equation}
or 
$$
\cD =   \{ (0,\delta ) \in \cD  : \frac{\Delta }{a} |
\delta \} \cup \cE _1' \cup \cE _2'  \, .  $$
where $\cE _1 $ and $\cE _2$ are a convenient or arbitrary partition
of those $\delta$ with $\frac{\Delta }{c} \not | \, 
\delta $. 

\textbf{Step 2.}
We now  define the entire functions
\begin{align}
 & \hspace{ 2 cm} \sigma_{\Lambda } (z) =   \label{defsig1} \\
&=  \sigma (p_1((AB)\inv z)) \prod _{\nu \in \cE _1} \Big(\beta _{p_1(B\inv \nu))  }\sigma
\Big)  \Big(p_1 \big( (AB)\inv z) \Big)
  \, \prod _{\nu \in \cE _2} \Big(\beta _{p_2(B\inv \nu )  }\sigma 
\Big)    \Big(p_2 \big( (AB)\inv z \big) \Big), \notag
  \end{align}
and 
\begin{align*}
 & \hspace{ 2 cm} \tilde \sigma_{\Lambda } (z)  =    \label{defsig2} \\
&=  \sigma (p_2((AB)\inv z)) \prod _{\nu \in \cE _1'} \Big(\beta _{p_1(B\inv \nu))  }\sigma
\Big)  \Big(p_1 \big( (AB)\inv z) \Big)
  \, \prod _{\nu \in \cE _2'} \Big(\beta _{p_2(B\inv \nu )  }\sigma 
\Big)    \Big(p_2 \big( (AB)\inv z \big) \Big). \notag
  \end{align*}


  \begin{prop}
    \label{mainproperty}
The  functions $\sigma _\Lambda$ and $ \tilde \sigma _\Lambda $   
 vanish on $\Lambda $. 
  \end{prop}

  \begin{proof}
    Let $k,l \in \mbZ $, $(0,\delta ) \in \cD $, $\nu = (0,\eta ) \in
    \cD $ and let   $\lambda = A\Big(B
\Big(\begin{smallmatrix}
  k \\ l 
\end{smallmatrix}\Big)
+
\Big(\begin{smallmatrix}
  0 \\ \delta 
\end{smallmatrix} \Big)\Big)$  be a general lattice point in $\Lambda
$. We write the coordinates  of the argument in \eqref{eq:d3}  as 
$$
p_1\big( (AB)\inv \lambda \big) = k  -c\delta /\Delta \qquad \text{
  and } \qquad p_2\big( (AB)\inv \lambda \big) = l+ a\delta / \Delta
\, .
$$
If we omit the normalizing factors of the Fock shifts, we have 
\begin{align*}
  \beta _{p_1(B\inv \nu )}\sigma (p_1((AB)\inv \lambda ) & \asymp
  \sigma (k-\frac{c\delta}{\Delta}+\frac{c\eta}{\Delta}) \, , \\
  \beta _{p_2(B\inv \nu )}\sigma (p_2((AB)\inv \lambda ) & \asymp
  \sigma (l+\frac{a\delta}{\Delta}-\frac{a\eta}{\Delta})  \, .
\end{align*}

If $(0,\delta ) \in \cE _0$, i.e., $\tfrac{\Delta}{c} | \delta $ and
thus 
$\delta = r \tfrac{\Delta}{c}$ for some $r\in \mbZ $, then 
$$
\sigma (p_1((AB)\inv \lambda )) \asymp  \sigma (k-c\delta  / \Delta )
= \sigma (k-r) = 0 \, . 
$$
So $\sigma $ vanishes on the cosets $(0,\delta ) + \Gamma $, whenever
$\tfrac{\Delta}{c} | \delta $. 

If $(0,\delta ) \in \cE _1 $, i.e., $\tfrac{\Delta}{c} \not
| \delta $, then the factor 
$$
  \beta _{p_1(B\inv \nu )}\sigma (p_1((AB)\inv \lambda )  \asymp
  \sigma (k-\frac{c\delta}{\Delta}+\frac{c\eta}{\Delta}) 
$$
vanishes for $\nu = (0,\delta ) \in \cE _1$. Likewise, if  $(0,\delta
)\in \cE _2$, then for $\nu \in \cE _2$ 
$$
  \beta _{p_2(B\inv \nu )}\sigma (p_2((AB)\inv \lambda )  \asymp
  \sigma (l+\frac{a\delta}{\Delta}-\frac{a\eta}{\Delta}) = 0 \, . 
$$
Thus every coset $(0,\delta )^T +\Gamma $ is annihilated by a single
factor of $\sigma _\Lambda $,  and  $\sigma  _\Lambda $  vanishes on
$\Lambda $.  

The proof for $\tilde \sigma _\Lambda $ is the same. 
  \end{proof}

\subsection{Interpolating Functions}

For the  interpolating function on $\Lambda $ we make a similar
ansatz. However, we need to pay special attention to the cosets in
$\cE _0$, where $\frac{\Delta}{c} | \delta $.  We partition $\cD  = \cE _0 \cup \cE _1 \cup \cE _2$ as in
\eqref{eq:vm1}. Evaluating $p_j\big( (AB)\inv \lambda \big), j=1,2$,
on these cosets, eventually leads to the following definitions.
Let $b_1, b_2$ be the basis vectors of $\Gamma $, namely
$b_1 = AB\big(\begin{smallmatrix}
  1 \\ 0 
\end{smallmatrix}\big)$ and $ b_2 = AB
\big(\begin{smallmatrix} 
  0 \\ 1
\end{smallmatrix}\big)$, and set
\begin{equation}
  \label{eq:n4}
b_3 = b_2 + A\Big(\begin{smallmatrix}
  0 \\ \Delta / c 
\end{smallmatrix}\Big) \in \bC ^2 \text{ and } \zeta = \frac{\langle
  b_3,b_2\rangle }{\|b_2\|^2} \in \bC \, .
\end{equation}
Recall that $\tau (w) = \sigma (w)/w, w\in \bC $ is interpolating for
$\mbZ $ and  define the entire function 
\begin{align*}
\tau_{\Lambda } (z) &=  \notag 
 \tau\big( p_1((AB)\inv z)\big)\,  \prod _{\nu \in \cE _1}
\Big(\beta _{p_1(B\inv \nu)  }\sigma \Big)  \Big(p_1 \big( (AB)\inv z\big) \Big) \\
 & \qquad \times \, \prod _{\nu \in \cE _2} \Big(\beta _{p_2(B\inv \nu )  }\sigma 
\Big)    \Big(p_2 \big( (AB)\inv z \big) \Big) 
   \prod _{r = c\delta /\Delta \in \mbZ \setminus
   \{0\}} \beta _{r\zeta } \sigma \Big( \frac{\langle z,
   b_2\rangle }{\|b_2\|^2} \Big)\, .
\end{align*}
The construction of $\tau _\Lambda $ is more subtle,  since we need to
include a finite number of extra factors $ \beta _{r\zeta } \sigma \Big( \frac{\langle z,
  b_2\rangle }{\|b_2\|^2}\Big)$.

\begin{prop} \label{interp1}
  $\tau _\Lambda $ is an interpolating function for $\Lambda $.
\end{prop}

\begin{proof}
The proof is similar to the proof of Proposition~\ref{mainproperty}.   Let $(0,\delta )
\in \cD $, $\nu = (0,\eta ) \in \cD $ and   $\lambda = A\big(B
\big(\begin{smallmatrix}
  k \\ l 
\end{smallmatrix}\big)
+
\big(\begin{smallmatrix}
  0 \\ \delta 
\end{smallmatrix} \big)\big)$  be a general lattice point in $\Lambda
$.  Then as before  
we have 
\begin{align*}
\tau\big (p_1((AB)\inv \lambda )\big) 
 &= \tau (k- \frac{c\delta}{\Delta}) \\
  \beta _{p_1(B\inv \nu )}\sigma (p_1((AB)\inv \lambda )) & \asymp
 \sigma (k-\frac{c\delta}{\Delta}+\frac{c\eta}{\Delta}) \\
  \beta _{p_2(B\inv \nu )}\sigma (p_2((AB)\inv \lambda )) & \asymp
  \sigma (l+ \frac{a\delta}{\Delta}-\frac{a\eta}{\Delta})  \, .\\
\end{align*}

If $\lambda = 0$ ($k=l=\delta =0$), then 
$$
\tau _\Lambda (0)  \asymp  \tau  (0)  \, \prod _{\eta \in
  p_1(\cE _1)} \sigma (\frac{c\eta }{\Delta}) \prod _{\eta \in
  p_2(\cE _2)} \sigma (\frac{-a\eta }{\Delta}) \neq 0
$$
since by definition $\tau  (0) = 1 $ and $ \frac{c\eta
}{\Delta}  \not \in \mbZ $, $\frac{-a\eta }{\Delta}\not \in \mbZ $
for $\nu \in \cE_1 \cup \cE_2$. 

If $(0,\delta ) \in \cE _1 \cup \cE
_2$,  then  
$\tau _\Lambda (\lambda
) = 0$ as in Proposition~\ref{mainproperty}. 

Finally,  if 
$(0,\delta ) \in \cE _0$,  then  $\delta  = r \Delta
 /c$ for some $r \in \mbZ $. Therefore 
$$
\tau\big( p_1((AB)\inv z)\big) = \tau (k- \frac{c\delta}{\Delta})  =
\tau (k-r) \, 
$$
and $\tau _\Lambda (\lambda ) = 0$, unless $k=r$. 

If $k=r$ or $\delta = r\Delta /c$, then, with the notation of
\eqref{eq:n4},  
$$ \lambda = 
 A\Big(B
\Big(\begin{smallmatrix}
  r \\ l 
\end{smallmatrix}\Big)
+
\Big(\begin{smallmatrix}
  0 \\ r\Delta/c  
\end{smallmatrix} \Big)\Big) = l AB \Big(\begin{smallmatrix}
  0 \\ 1
\end{smallmatrix} \Big) + r \Big(AB \Big(\begin{smallmatrix}
  1 \\  0  
\end{smallmatrix} \Big) + A \Big( \begin{smallmatrix}
  0 \\ \Delta/c  
\end{smallmatrix} \Big) \Big) = l b_2 + rb_3 \, .
$$

We consider the factor $ \beta _{r\zeta } \sigma \Big( \frac{\langle z,
   b_2\rangle }{\|b_2\|^2}\Big)$ of $\tau _\Lambda $ with $ \zeta = \frac{\langle
  b_3,b_2\rangle }{\|b_2\|^2}$  and evaluate at
 $\lambda $: 
\begin{align*}
\beta _{r\zeta } \sigma \big( \frac{\langle \lambda ,
  b_2\rangle}{\|b_2\|^2}\big) & \asymp \sigma \big(    \frac{\langle \lambda ,
  b_2\rangle}{\|b_2\|^2} - r\zeta \big) \\
&= \sigma \big(    \frac{\langle l b_2 + rb_3, b_2\rangle }{\|b_2\|^2}
- r \frac{\langle b_3,b_2\rangle}{\|b_2\|^2} \big) = \sigma (l) = 0 \, .
\end{align*}
Altogether we have shown that $\tau _\Lambda (\lambda ) = \tau
_\lambda (0) \delta _{\lambda ,0}$ and that $\tau _\Lambda (0) \neq
0$, so that $\tau _\Lambda $ is an interpolating function for $\Lambda
$.




  \end{proof}
\section{Failure of  Sampling in  Fock Space}

 We now apply the construction of sigma-type functions to prove the
 failure of sampling for certain lattices of density $>1$. Such examples appear naturally 
 if one considers a sigma-type function of sufficiently small growth. 
 Its zero set is a "lattice" of hyperplanes, and  one can then choose
 a discrete lattice of arbitrarily large  density which belongs to these
 hyperplanes. For instance, $\sigma _0(z_1,z_2) = \sigma (z_1) \sigma
 (z_2)$ is constructed to vanish on $\mbZ ^2$, but its zero set is the
 union of the complex lines $\{k\} \times \bC $ and $\bC \times \{l\}$
 for $k,l \in \mbZ $. Consequently $\sigma _0$ vanishes on every
 lattice $\epsilon \mbZ \times  \mbZ $, which has  density $\epsilon ^{-2}>>1$.  
 
  Surprisingly this is not the only possibility of building such examples.

\begin{tm} \label{fail}
  Let $q=q_1 + i q_2 \in \mbZ $ with $|q|\geq 2$,  $\gamma ^2 + 1/|q|^2=1$ and $$\Lambda =
  \begin{pmatrix}
    1 & \frac{1}{q} \\
 0 & \gamma 
  \end{pmatrix} \mbZ ^2 \, .
$$
Then $\Lambda $ fails to be a set of sampling for $\cF $. 
\end{tm}

\begin{proof}
  According to Proposition~\ref{chargauss} it suffices  to find an $F\in
  \cF ^\infty _2 $,  
  such that $F(\lambda ) = 0$ for all $\lambda \in \Lambda $. 
 We  apply Proposition~\ref{mainproperty} to the  sublattice $\Gamma = AB\mbZ^2$ of $\Lambda $ defined
by the matrix  
$$
B =
\begin{pmatrix}
  1 & -\bar{q} \\ 0 & |q|^2
\end{pmatrix} \,  
$$
with inverse 
$
B\inv =
\left ( 
\begin{smallmatrix}
  1 & 1/q \\ 0 & 1/|q|^2 
\end{smallmatrix}
\right )
$. 
Then $AB = \mathrm{diag}\, (1, \gamma |q|^2)$ and thus 
                 $\Gamma = AB (\mbZ )^2 = \mbZ \times \gamma |q|^2
                 (\mbZ )$  
 possesses  an orthogonal
basis.  Then $\Delta = \det B = |q|^2$, and 
by Lemma~\ref{l3} the cosets of $\mbZ ^2 / B\mbZ ^2$ are represented by the
set $\cD  = \{ (0, \delta _1 + i \delta _2) \in \mbZ ^2: 0\leq \delta
_1, \delta _2 < 
|q|^2\}$. 

Consequently the sigma-type  function of $\Gamma $ is 
$$
\sigma _\Gamma (z)  = (\sigma \otimes \sigma )\big( (AB)\inv z
\big) = \sigma (z_1) \sigma (\frac{z_2}{\gamma |q|^2})
\, .
$$
Following the recipe of \eqref{eq:vm1} and \eqref{defsig1}   we  partition  
the  coset representatives $\cD  $ into 
$$
\cE_0 = \{ (0, \delta ) \in \cD :  q |  \delta \} = \{ (0,\delta ):
\delta = q (m+in), m,n \ = 0, \dots , |q| -1 \} \, ,
$$
and $\cE _1 = \emptyset $, and     
 $\cE _2 = \cD \setminus \cE _0 =  \{ (0, \delta ) \in \cD :  q \not |
 \,  \delta \}
 $. 
 
Then $\mathrm{card}\, \cE _0  =  |q|^2$ and $\mathrm{card}\, \cE _2  =
|q|^4 - |q|^2$.  In the notation of Section~3.2 $p_2((AB)\inv z) =
z_2/(\gamma |q|^2)$ and $p_2(B\inv \nu ) = \delta /|q|^2$. Then
\eqref{defsig1} yields the explicit formula 
\begin{align}
\sigma _\Lambda (z_1,z_2) &=  \sigma (z_1) \prod
_{\delta \in \cE _2} (\beta _{\delta / |q|^2} \sigma )
\big(\frac{z_2}{\gamma |q|^2}\big) \notag \\
&= \sigma (z_1) \, \prod
_{\delta \in \cE _2} e^{\pi \bar{\delta } z_2 /|q|^2} \sigma
\big(\frac{z_2}{\gamma |q|^2}- 
\frac{\delta}{|q|^2}\big) \, e^{-\frac{\pi}{2} \frac{|\delta
    |^2}{|q|^4}} \,  \label{vn1} \,
\end{align}
By Proposition~\ref{mainproperty}, 
$\sigma _\Lambda $ vanishes on $\Lambda $.   




Finally we need to   check  the growth of $\sigma _\Lambda $.   Since $\beta _{w_0}  $  is an isometry on $\cF
^\infty _1$ by \eqref{eq:d9bbb}, and since there are exactly
$|q|^4-|q|^2$ factors  in the product over $\cE _2$, we obtain 
\begin{align*}
|\sigma _\Lambda (z_1,z_2)| & = |\sigma (z_1)| \prod
_{\delta \in \cE _2} |(\beta _{\delta / |q|^2} \sigma )
\big(\frac{z_2}{\gamma |q|^2} \big) | \\
& \leq \exp \big(\frac{\pi }{2}
|z_1|^2 \big) \prod _{\delta \in \cE _2} \exp \big( \frac{\pi}{2}
\frac{|z_2|^2}{\gamma ^2|q|^4} \big) \\
& =  \exp \big(\frac{\pi }{2}
(|z_1|^2 + \frac{|q|^4-|q|^2}{\gamma ^2 |q|^4} |z_2|^2 ) \big) \\
& =  \exp \big(\frac{\pi }{2}
(|z_1|^2 + |z_2|^2\big) \, ,
  \end{align*}
where in the last identity we have used $1-\tfrac{1}{|q|^2} = \gamma
^2$. 
Thus $\sigma _\Lambda \in \cF ^\infty _2$ and by
Proposition~\ref{chargauss} $\Lambda $ cannot be a set of sampling for
$\cF $. 
\end{proof}

\begin{cor} \label{cormany}
Let $\Lambda ' $ be a lattice of the form
$$
\Lambda '=
\begin{pmatrix}
  \alpha  &0 \\ 0 & \beta 
\end{pmatrix}
\begin{pmatrix}
  1 & \frac{1}{q} \\ 0 & \gamma
\end{pmatrix} \mbZ^2
$$
with $\alpha ,\beta \geq 1 $ and $\gamma ^2 + \tfrac{1}{q^2}=
1$. Then $\Lambda ' $ is not a set of sampling for $\cF $.   
\end{cor}

\begin{proof}
Let $D = \mathrm{diag} (\alpha, \beta )$, then $\Lambda '= D \Lambda $
with $\Lambda $ as in Theorem~\ref{fail}. 
  
 According to Theorem~\ref{fail}  there exists a non-zero  function $F \in
\cF ^\infty_2 $, such that $F(\lambda ) = 0$ for all $\lambda \in
\Lambda $. Set $\tilde{F}(z) = F(D\inv z)$.
If  $\lambda ' \in \Lambda '$, i.e. $\lambda '  = D\lambda $ for
some $\lambda \in \Lambda $, we have $\tilde F (\lambda ') = F(D\inv D
\lambda ) = F(\lambda ) = 0$ and 
$$
|\tilde F (z) | = |F(D\inv  z)| \leq C e^{\frac{\pi }{2} |D\inv z|^2}
\leq C e^{\frac{\pi }{2} | z|^2} \, .
$$
So $\tilde F \in \cF ^\infty _2 $ and $\tilde F $ vanishes on $\Lambda
'$. Therefore Proposition~\ref{chargauss} implies that  $\Lambda '$
fails to be a set of sampling for $\cF $.  
\end{proof}

\begin{cor} \label{corun}
  Let $p,q\in \bN$, $q\geq 2$,  $\gamma ^2 + 1/q^2=1$ and $$\Lambda =
  \begin{pmatrix}
    1 & \frac{p}{q} \\
 0 & \gamma 
  \end{pmatrix} \mbZ^2 \, .
$$
Then $\Lambda $ fails to be a set of sampling for $\cF $.   
\end{cor}

\begin{proof}
  The proof is almost  the same as of Theorem~\ref{fail}. We  choose
  the sublattice $\Gamma $ determined by the matrix
$$B=
\begin{pmatrix}
  1 & -p \\ 0 & q
\end{pmatrix} \, .
$$
Then  $\Gamma = AB \mbZ ^2 = \Big(
\begin{smallmatrix}
  1 & 0 \\ 0 & \gamma q
\end{smallmatrix}\Big) \mbZ ^2$ has an orthogonal basis, and  $B\inv  = \Big(
\begin{smallmatrix}
  1 & \tfrac{p}{q} \\ 0 & \tfrac{1}{q} 
\end{smallmatrix}\Big)$ and $\cD = \{ (0,\delta _1+i\delta _2): 0\leq
\delta _j <q \}$. The sigma function is 
$$
\sigma _\Lambda (z_1,z_2) = \sigma (z_1) \prod _{\eta \in \cD
  \setminus \{ 0\}} \beta _{\eta /q} \sigma \big(\frac{z_2}{\gamma
q}\big)
$$
with growth 
$$
|\sigma _\Lambda (z) | \leq e^{\tfrac{\pi }{2} |z_1|^2} \,
e^{\tfrac{\pi }{2} (q^2-1) \frac{|z_2|^2}{\gamma ^2q^2}} =
e^{\tfrac{\pi }{2} |z|^2} \, ,$$
 because $1  - q^{-2} = \gamma ^2$. 
\end{proof}

One may wonder how we chose the sublattice $\Gamma $. Our guiding
principle was to find a sublattice with an orthogonal basis because in
this case the norm $\|(AB)\inv \|_{\mathrm{op}}$ is minimized.  We
do not know  how to make Theorem~\ref{fail} work in greater
generality.

Consider the one-parameter family of matrices $A_t=\big(
\begin{smallmatrix}
  1 & \frac{2}{5} \\ 0 & t 
\end{smallmatrix}\big) $ with $t^2+(2/5)^2\geq 1$ (as we assume a
reduced basis).  Corollary~\ref{corun} and \ref{cormany} say
that the lattices $\Lambda _t
= A _t \mbZ ^2$  fail to be  sampling  for $t \geq  \sqrt{24}/5$, but
we do not know what happens for $\sqrt{21}/5\leq   t <  \sqrt{24}/5$.

\section{A Weak  Sampling Formula}

Theorem~\ref{fail} and Corollaries~\ref{cormany} and~\ref{corun} show
that many lattices 
with density $>1$ fail to be sampling sets for $\cF _2^2$.
By contrast, we have not succeeded to use the interpolating functions
of Proposition~\ref{interp1} to prove positive results about sampling
lattices.  
Nevertheless, we can prove a slightly weaker sampling theorem which holds even for dimensions $d>2$.

This section is a symbiosis of complex analysis (construction of
interpolating functions) and Gabor analysis (duality theory of Gabor
frames). In fact, we translate a weak reconstruction formula of
Feichtinger and Zimmermann~\cite{FZ98} into a Lagrange-type
reconstruction for entire functions.

For the discussion we need some more background from Gabor
analysis. The modulation space $M^1(\rd )$ is the subspace  of $\lrd $
for  which  the norm 
$$\|h\|_{M^1} = \int _{\rdd}  |\langle h, \pi _z \phi \rangle |\, dz$$
is finite, where   $\phi $ is the normalized Gaussian. 
Its dual is the space of tempered distributions such
that $\sup _{z\in \rdd } |\langle h, \pi _z \phi \rangle| < \infty $. Then
 $(M^1)^* = M^{\infty }$, and the duality  is given  via the Bargmann
transform as 
$$
\langle h,k\rangle _{M^1 \times M^{\infty }} = \int_{\cd } Bh(z)
\overline{Bk(z)} \, e^{-\pi |z|^2} \, dz  = \langle Bh, Bk\rangle \, .
$$
  Introducing the  Fock
space  $\cF ^1_d$ consisting of all entire functions with finite norm 
$$
\|F\|_{\cF ^1_d} = \int _{\cd } |F(z)| \, e^{-\pi  |z|^2/2} \,
dz  <\infty \, , 
$$
we can identify  $M^1$ and $M^\infty $ as  the pre-images of $\cF ^1_d$
and $\cF ^\infty _d$ of the Bargmann transform~\cite{G2}.

The detailed analysis of the duality theory of Gabor frames led
Feichtinger and Zimmermann~\cite{FZ98} to  the theory of
weak dual pairs.  We will
apply the following version of the duality
theory~\cite[Thm.~3.5.12]{FZ98}. 

\begin{prop}[Weak duality] \label{weakfz}
Let $\Lambda \subseteq \rdd $ be a
lattice with adjoint lattice $\Lambda ^\circ$,  $g\in M^1(\rd ) $ and $\gamma \in
M^\infty (\rd )$. Then the following are equivalent:

(i) Biorthogonality on the adjoint lattice: 
\begin{equation}
  \label{eq:vm8}
  \tfrac{1}{\mathrm{vol}\, (\Lambda )} \langle \gamma , \pi_\mu
g\rangle = \delta _{\mu ,0} \text{ for all } \mu \in \Lambda ^\circ  \, .
\end{equation}

(ii) For every $f,h\in M^1(\rd )$ we have 
\begin{equation}
  \label{eq:vn46}
  \langle f,h\rangle = \sum _{\lambda \in \Lambda } \langle f, \pi
  _\lambda g \rangle \langle \pi _\lambda \gamma , h\rangle
\end{equation}
with absolute convergence of the sum. 
\end{prop}
The identity  \eqref{eq:vn46}
can be interpreted as a reconstruction formula
$$
f = \sum _{\lambda \in \Lambda } \langle f, \pi
  _\lambda g \rangle \pi _\lambda \gamma
$$
for $f\in M^1$, but with  convergence in the weak$^*$-topology  on $M^\infty$.

By applying the Bargmann transform,  Proposition~\ref{weakfz} with
$g(x) =  2^{d/4} e^{-\pi |x|^2}$ is translated  into 
 the following Lagrange interpolation formula for $\cF ^1_d$. 

 \begin{cor} \label{dualentire}
Let $\Lambda \subseteq \cd $ be a complex 
lattice with adjoint lattice $\Lambda ^\circ \subseteq \cd$  and $\tau
\in
\cF ^\infty _d$. Then the following are equivalent:

(i) $\tau $ is interpolating on $\duall $.

(ii) For all $F\in \cF ^1_d$
\begin{equation}
  \label{eq:m8}
  F(z) = \sum _{\lambda \in \Lambda } F(\lambda ) e^{\pi \bar{\lambda
    } \cdot z} \tau(z-\lambda ) \, e^{-\pi |\lambda |^2} \, . 
\end{equation}
The series expansion converges weakly in the sense that for all $H\in
\cF ^1_d$
$$
\int _{\cd }  F(z) \overline{H(z)} e^{-\pi |z|^2}  \, dz = \sum _{\lambda \in
  \Lambda } F(\lambda ) e^{-\pi |\lambda |^2} \,  \int _{\cd } e^{\pi \bar{\lambda
    } \cdot z} \tau(z-\lambda )   \overline{H(z)} e^{-\pi |z|^2}  \,
  dz \, .
  $$
 \end{cor}

 \begin{proof}
 This follows from the properties of the Bargmann transform: $\langle
 f,h \rangle = \int _{\cd } Bf(z)  \overline{Bh(z)} e^{-\pi |z|^2}  \,
 dz $, and $\langle f, \pi _\lambda \phi \rangle  = Bf(\lambda )
 e^{-\pi |\lambda |^2/2}$ and the fact that the Bargmann transform is
 an isomorphism from $M^1$ \emph{onto} $\cF ^1_d$ and from $M^\infty $
 onto $\cF ^\infty _d$.  
 \end{proof}

 Combining Corollary~\ref{dualentire} with the construction of interpolating functions we 
 obtain the following Lagrange interpolation formula. 
\begin{tm} \label{weaksamp}
Let $\Lambda = S \mbZ ^d$ where $S$ is  an upper triangular matrix
with diagonal $(\gamma _1, \dots , \gamma _d)$ as in \eqref{eq:r1}.
Assume that $\max _{j=1, \dots , d} \gamma _j = 1$.

(i) Then there exists $\gamma \in M^\infty $ such that 
$\tfrac{1}{\mathrm{vol}\, (\Lambda )} \langle \gamma , \pi_\mu
g\rangle = \delta _{\mu ,0}$  for all $ \mu \in \Lambda ^\circ $ and
the weak reconstruction formula \eqref{eq:vn46} holds.

(ii) Equivalently, there exists an interpolating function  $\tau \in
\cF ^\infty _d$ for
the adjoint lattice $\duall $,  such that the weak Lagrange
interpolation formula \eqref{eq:m8} holds.  In particular, $\Lambda $
is a set of uniqueness for $\cF ^1_d$. 
\end{tm}

\begin{proof}
  The adjoint lattice of $\Lambda $  is $(S\inv )^* \mbZ ^d$, and the diagonal of
  $(S\inv )^*$ is $(\gamma _1\inv , \dots , \gamma _d \inv )$. Then
  the entire function
  $$
\tau (z) = \tau (z_1,z_2, \dots , z_d)= \prod _{j=1}^d \frac {\sigma(\gamma_j z_j)}{z_j}
$$
is interpolating for $\duall $. This was already proved in
\cite{G1}. Furthermore, by the growth estimate for the Weierstrass
sigma-function \eqref{hayman} we have
$\big|\frac{\sigma(\gamma_j z_j)}{z_j}\big| \leq C e^{\pi \gamma _j^2
  |z_j|^2 / 2}$. Consequently,
$$
|\tau (z) | \leq C' e^{\pi (\max \gamma _j^2 )\, |z|^2/2} \, .
$$
Consequently, $\tau \in \cF ^\infty _d$. However, since $\gamma _{l}
= 1$ for some $l$, we have $\tau \not \in \cF ^2_d$.
We now choose $\gamma \in M^\infty $, so that $B\gamma = \tau $. The
statement now follows from the assertion of Proposition~\ref{weakfz}
and Corollary~\ref{dualentire}. 
\end{proof}

Obviously our results are far from complete and  should be considered  a
 collection of expected and of  surprising  examples. At this time we do
 not even understand the sampling property of the class of lattices
 $\Lambda = A \mbZ ^2$ with $A=\big(
 \begin{smallmatrix}
   1 & \beta \\ 0 & \gamma 
 \end{smallmatrix}\big) $ and $|\beta |^2 + \gamma ^2 = 1$. As we have seen,
 for certain
 values of $\beta $, $\beta = 1/q, q\in \mbZ , |q|\geq 2$,  the lattice
 $\Lambda $ is not sampling, but nothing else is known.


 \bibliographystyle{abbrv}
 \bibliography{general,new}

\def\cprime{$'$} \def\cprime{$'$} \def\cprime{$'$} \def\cprime{$'$}
  \def\cprime{$'$} \def\cprime{$'$}

\end{document}